ames invested $1,000 in large U.S. stocks at the beginning of 2012. This investment earned 15.00 percent in 2012, 30.00 percent in 2013, 11.00 percent in 2014, and 1.00 percent in 2015. What return did he earn in the average year during the 2012–2015 period? (Round answer to 2 decimal places, e.g. 1.52.)

Excel Template
(Note: This template includes the problem statement as it appears in your textbook. The problem assigned to you here may have different values. When using this template, copy the problem statement from this screen for easy reference to the values you’ve been given here, and be sure to update any values that may have been pre-entered in the template based on the textbook version of the problem.)

Return earned in the average year

%
Click if you would like to Show Work for this question:
Answer #1

solution : 20%+ 30.00%t 11.00%) +1.00 1. = 57/4 = 14.25 Hence the Cossect anwes is 14.966% * please date *